OpenStudy (anonymous):

f(x)=x^2+4x+1, identify the vertex and axis of symmetry.

OpenStudy (anonymous):

complete the square to get \[f(x)=(x+2)^2+1-4=(x+2)^2-3\]

OpenStudy (anonymous):

vertex is (-2,-3) and axis of symmetry is x=-2

OpenStudy (anonymous):

or you could just use \[-\frac{b}{2a}=-\frac{4}{2}=-2\] and \[f(-2)=(-2)^2+4\times -2+1=4-8+1=-3\]

OpenStudy (anonymous):

thanks A LOT

OpenStudy (anonymous):

\[-\frac{b}{2a}\] is usually easier because you do not have to keep track of adding and subtracting. to find the y coordinate just substitute